OpenStudy (anonymous):

If you help me out, I will give you 2 medals and I will fan you for your hard work. I have 2 accounts so I can give you 2 medals, my questions are below and it is 7th grade math. If I just posted this, the math questions may not be there.

OpenStudy (anonymous):

1. Solve the equation. –9v – 5 = –95 (1 point) A. 17 B. 11 C. 10 D. –10 2. Solve the equation. – 10 = –12 (1 point) A. –66 B. 6 C. –6 D. 66 3. Solve the equation. + 6 = 10 (1 point) A. 44 B. 30 C. 20 D. –20 4. Solve the equation. 3(1 – 2k) = –5k (1 point) A. –1 B. 1 C. 2 D. 3 5. Solve the equation. 3n + 2= 8 + 2n (1 point) A. 3 B. 4 C. 5 D. 6 6. Simplify the expression. –9 – 6(w + 5) (1 point) A. –15w + 5 B. –15w + 30 C. –6w – 39 D.–6w + 21 7. Solve the equation. 4(y – 4) = 8 (1 point) A. –2 B. 2 C. 4 D. 6 For questions 8–10, solve the inequality. 8. c + 6 < –20 (1 point) A. c < –26 B. c < –14 C. c < 26 D. c < 14 11. \[\frac{ P }{ 8 }\]≥ –5 (1 point) P\8 belongs in the sentence, it messed up. A. p ≥ 40 B. p ≥ 3 C. p ≥ –13 D. p ≥ –40 10. –5p > –30 (1 point) A. p > 6 B. p < 6 C. p > –35 D. p < 35 13. Describe the number of solutions for the equation. 10(x – 2) = 10x (1 point) A. one solution B. no solution C. infinite solutions 14. Describe the number of solutions for the equation. –3(x – 5) = 3x + 9 (1 point) A. one solution B. no solution C. infinite solutions 15. Solve the equation. 5x + 10 = 15 (1 point) A. 3 B. 1 C. 5 D. 25 Thank you very much in advance.

OpenStudy (anonymous):

i know #15.)

OpenStudy (anonymous):

first u do 15-10 then u get 5 then u do 5 divided by 5 which gives u....

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

it is B. 1

OpenStudy (anonymous):

15=b

OpenStudy (anonymous):

sorry i am not going to do all of them but i will try to do 2 or 3

OpenStudy (anonymous):

That helps, but if you do all of them I'll give you 2 medals and I will fan you as promised. And thankyou Pig president. Sorry I'm just really failing math and need help ASAP

OpenStudy (anonymous):

i think i know #7.

OpenStudy (anonymous):

7. Solve the equation. 4(y – 4) = 8 ok so.. hold on

OpenStudy (anonymous):